OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of 2 over 3? 2x + 3y = 3 2x – 3y = 3 3x + 2y = 3 3x – 2y = 3

OpenStudy (anonymous):

y = mx+b 3 = m(6) + b 3 = (2/3)(6) + b 3 = 4 + b b = -1 y = (2/3)x -1 3y = 2x - 3 2x - 3y -3 = 0 2x - 3y = 3